OpenStudy (anonymous):

Use the discriminant to find the number of real solutions of the equation 3x^2 – 5x + 4 = 0. Show your work.

OpenStudy (abdullahm):

The discriminant of an equation of the form \(\Large\bf ax^2 + bx + c \) is \(\Large\bf b^2 - 4ac\)

OpenStudy (aaronq):

Using what the person above provided, \(\sf D=(5)^2-(4*3*4)\) D=25-54 D<0 look at this table for the summary of what D means

OpenStudy (astrophysics):

@ParthKohli